# On the Role of Context in Reading Time Prediction

### v2

@@ 1 Introduction
This raises the question of whether there is a more suitable way to estimate the effect that context has on reading time.author=Andreas,color=blue!40,size=,fancyline,caption=,]Maybe we should also mention that they are correlated with frequency before saying this We argue that, given that frequency is known to play an important role in processing effort (Broadbent, 1967; Inhoff and Rayner, 1986; Rayner and Duffy, 1986; Bybee, 2006), a more interesting construct to analyze should be what context contributes beyond what is already captured by frequency. To obtain a predictor that represents just that, we propose a technique where we project surprisal onto the orthogonal complement of frequency, ensuring that they are uncorrelated. This process effectively disentangles the contextual and non-contextual information into different covariates in our regressions and closely resembles residualization. [^1] See App. D for more discussion on residualization.
@@ 3 Disentangling the Effect of Context
Thus, if ${\mathbf{Y}_{\mathrm{H}}}$ and ${\mathbf{I}_{\mathrm{H}}}$ are mean-centered the covariance between ${\mathbf{Y}_{\mathrm{H}}}$ and $\mathrm{proj}_{{\mathbf{Y}^{\perp}_{\mathrm{H}}}}({\mathbf{I}_{\mathrm{H}}})$ will be 00, i.e., they will be decorrelated.author=Andreas,color=blue!40,size=,fancyline,caption=,]Ryan suspects that the coefficeints will be diffeerent if we dont mean center The random variable $\mathrm{proj}_{{\mathbf{Y}^{\perp}_{\mathrm{H}}}}({\mathbf{I}_{\mathrm{H}}})$ constitutes a new predictor variable, which we term orthogonalized surprisal. In words, orthogonalized surprisal represents the effect of context and is disentangled from frequency. [^7] The method presented can in principle be applied to any pair of predictor variables that live in ${\mathcal{H}}$, e.g., taking ${\mathbf{M}_{\mathrm{H}}}$ instead of ${\mathbf{I}_{\mathrm{H}}}$ yields an equivalent context predictor. We also consider orthogonalized unit length in our experiments in § 4. author=Andreas,color=blue!40,size=,fancyline,caption=,]Low priority but nice to have: Prove that doing this with PMI instead would be equivalent

### v3

@@ 1 Introduction
This raises the question of whether there is a more suitable way to estimate the effect that context has on reading time. We argue that, given that frequency is known to play an important role in processing effort (Broadbent, 1967; Inhoff and Rayner, 1986; Rayner and Duffy, 1986; Bybee, 2006), a more interesting construct to analyze should be what context contributes beyond what is already captured by frequency. To obtain a predictor that represents just that, we propose a technique where we project surprisal onto the orthogonal complement of frequency, ensuring that they are uncorrelated. This process effectively disentangles the contextual and non-contextual information into different covariates in our regressions and closely resembles residualization. [^1] See App. D for more discussion on residualization.
@@ 3 Disentangling the Effect of Context
Thus, if ${\mathbf{Y}_{\mathrm{H}}}$ and ${\mathbf{I}_{\mathrm{H}}}$ are mean-centered the covariance between ${\mathbf{Y}_{\mathrm{H}}}$ and $\mathrm{proj}_{{\mathbf{Y}^{\perp}_{\mathrm{H}}}}({\mathbf{I}_{\mathrm{H}}})$ will be 00, i.e., they will be decorrelated. The random variable $\mathrm{proj}_{{\mathbf{Y}^{\perp}_{\mathrm{H}}}}({\mathbf{I}_{\mathrm{H}}})$ constitutes a new predictor variable, which we term orthogonalized surprisal. In words, orthogonalized surprisal represents the effect of context and is disentangled from frequency. [^7] The method presented can in principle be applied to any pair of predictor variables that live in ${\mathcal{H}}$, e.g., taking ${\mathbf{M}_{\mathrm{H}}}$ instead of ${\mathbf{I}_{\mathrm{H}}}$ yields an equivalent context predictor. We also consider orthogonalized unit length in our experiments in § 4.